# Multi-finger Manipulation via Trajectory Optimization with Differentiable Rolling and Geometric Constraints

### v1

@@ IV-A3 Wrench Balance Constraint
We introduce wrench balance as the dynamics constraint for our quasi-static system. We use the direct collocation method to frame the system dynamics by imposing dynamics constraints $f_{dynamics}(\mathbf{s}_{t},\mathbf{u}_{t},\mathbf{s}_{t+1})=0$ on each collocation point. The collocation method has been widely used in trajectory optimization to handle complex dynamics [27, 28, 29], and often has better numerical properties than shooting methods.
@@
$-m\mathbf{g}+\sum_{i}\mathbf{f}_{i,t}+\mathbf{f}_{e,t}=0$ (5)
$\begin{split}&-m\mathbf{r_{com}}(\bm{\theta}_{t+1})\times\mathbf{g}+\sum_{i}\mathbf{f}_{i,t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})\\ &+\mathbf{f}_{e,t}\times\mathbf{r}(\mathbf{c}_{e,t},\mathbf{o}_{t})=0,\end{split}$ (6)
where $K_{p}$ is the gain matrix for the PD controller, $\bm{\tau}_{g}(\mathbf{q}_{t})$ is the torque required to counteract the gravity at each joint, and $\mathbf{g}=[0,0,9.8]^{T}$ is the gravity vector. $\mathbf{r_{com}}$ is the vector represented in the world frame pointing from the object frame origin to its CoM. $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ are the force that the $i$th finger and the environment apply to the object respectively. The visualization of the variables used in the force balance is shown in Fig. 3. Eq. 4 refers to the torque balance for robot joints. Eq. 5 and Eq. 6 refer to the force and torque balance for the object respectively.
Having both control input $\hat{\mathbf{u}}_{t}$ and contact forces $\mathbf{f}_{i,t}$ might seem redundant, since one can possibly derive $\mathbf{f}_{i,t}$ from $\mathbf{s}_{t}$ and $\hat{\mathbf{u}_{t}}$. However, explicitly computing the contact forces is not trivial [30]. Previous work has proposed to include computing contact forces as part of the optimization problem [31, 29]. Similar to those methods, we add the forces $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ as decision variables and compute them within the optimization. The wrench balances also ensure the consistency of control actions $\hat{\mathbf{u}}_{t}$ and forces.
@@ IV-A4 Friction Constraints
In this work, we use the Coulomb friction constraint: $||\mathbf{f}_{i,t}^{t}||\leq\mu||\mathbf{f}_{i,t}^{n}||,$ where $\mathbf{f}_{i,t}^{t}$ is the tangential force, $\mathbf{f}_{i,t}^{n}$ is the normal force, and $\mu$ is the friction coefficient. However, satisfying the second-order friction constraint can be numerically unstable when the tangential force is close to 0. To address this, we use a linearized 4-sided friction cone [32], formulated as:
$A(\mathbf{n}_{i,t},\mu)\mathbf{f}_{i,t}\leq 0,$ (7)
where the matrix $A$ depends on the contact normal $\mathbf{n}_{i,t}$ and the friction coefficient $\mu$.
@@ IV-A5 Trajectory Optimization
Our trajectory optimization problem is:
$\min_{\begin{subarray}{c}\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{T};\\ \mathbf{u}_{1},\mathbf{u}_{1},\cdots,\mathbf{u}_{T}\end{subarray}}J_{goal}(\bm{\tau})+J_{smooth}(\bm{\tau})$ (8)
@@
$-m\mathbf{g}+\sum_{i}\mathbf{f}_{i,t}+\mathbf{f}_{e,t}=0$ (14)
$\begin{split}&-m\mathbf{r_{com}}(\bm{\theta}_{t+1})\times\mathbf{g}+\sum_{i}\mathbf{f}_{i,t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})\\ &+\mathbf{f}_{e,t}\times\mathbf{r}(\mathbf{c}_{e,t},\mathbf{o}_{t})=0\end{split}$ (15)
$A(\mathbf{n}_{i}(\mathbf{s}_{t}),\mu)\mathbf{f}_{i,t}\leq 0.$ (16)
Eq. 11, 12, and 16 refer to $f_{contact}$, $f_{kinematics}$, and $f_{friction}$, and Eq. 13, 14, and 15 refers to $f_{balance}$.
@@ IV-B Geometry Parametrization
As described above, we utilize geometry information to formulate the constraints, such as contact constraints, and kinematics constraints. However, it is generally intractable to take geometry information such as meshes as input, i.e., non-primitive shapes such as robot fingers cannot be described with just a few parameters (e.g., radius, length, height). It is even more challenging to compute geometry-related gradients for use in trajectory optimization algorithms.
In this section, we describe how we use a differentiable SDF with a sampling-based method to estimate geometry-related variables and constraints. Those include the distance function $\Phi_{i}(\mathbf{q}_{t},\mathbf{o}_{t})$ for contact constraints, contact points $\mathbf{c}_{i,t}$, the jacobian matrix at the contact point $J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})$, and the contact normal $\mathbf{n}_{i,t}$ for kinematics constraints. $\mathbf{n}_{i,t}$ is also used in formulating the friction constraints. Gradients of the constraints above w.r.t. the state $\mathbf{s_{t}}$ and actions $\mathbf{u}_{t}$ are also very important since gradients describe how to change the states and actions to satisfy constraints in a lower-dimensional manifold. In this section, we will ignore the timestep subscript $t$ in our equations for simplicity.
We assume the manipulated object and the robot are described by primitive shapes and triangle meshes respectively. Primitive shapes include boxes, cylinders, spheres, and their combinations. Many objects in real-world applications can be simplified as primitives or their combinations, e.g., a screwdriver as a combination of two cylinders with different radii. Primitives have an analytical formulation of their geometry and SDF, enabling efficient computation of contacts and distances. In the following text, we will utilize the function $\phi_{k}(\mathbf{p}_{j},\mathbf{o})$ to query the SDF of the $k$th primitive of configuration $\mathbf{o}$ at $\mathbf{p}_{j}$. The subscript $k$ is introduced because, in practice, the object can be a combination of multiple primitives. Therefore, the distance between the point $\mathbf{p}_{j}$ and the object can be written as $\phi(\mathbf{p}_{j},\mathbf{o})=\min_{k}\phi_{k}(\mathbf{p}_{j},\mathbf{o})$
@@
With the assumption of the primitive objects, we have access to the SDF of the object: $\phi(\mathbf{p}_{j},\mathbf{o})$. The distance between the $i$th finger and the object can be written as:
$\Phi_{i}(\mathbf{q},\mathbf{o})=\min_{j\in P_{i}}\phi(\mathbf{p}_{j},\mathbf{o}),$ (17)
where $P_{i}$ is the set of points sampled on the surface of the $i$th fingertip. While this provides an accurate estimation with enough sampled points, the minimization operation is not differentiable and also throws away information from the other points. To address the differentiablity problem, we use the softmin to produce a weighted summation over all the points. The weight $h_{j}$ for the $j$th point is given by:
$h_{j}=\frac{\exp(-\delta\phi(\mathbf{p}_{j},\mathbf{o}))}{\sum_{k}\exp(-\delta\phi(\mathbf{p}_{k},\mathbf{o}))},$ (18)
where $\delta$ is the temperature. In practice, we use a high temperature since we want the softmin approximation to be as close as possible to the actual distance. The estimated distance is written as:
$\Phi_{i}(\mathbf{q},\mathbf{o})=\sum_{j\in P_{i}}h_{j}\phi(\mathbf{p}_{j},\mathbf{o}).$ (19)
@@
$\mathbf{c}_{i}=\sum_{j\in P_{i}}h_{j}\mathbf{p}_{j}.$ (20)
$J_{\mathbf{c}_{i}}(\mathbf{q})=\sum_{j\in P_{i}}h_{j}J_{\mathbf{p}_{j}}(\mathbf{q}).$ (21)
Additionally, we also need to estimate the contact normal $\mathbf{n}_{i}(\mathbf{p},\mathbf{o})$ w.r.t. the object surface. The contact normal at the point $\mathbf{p}_{j}$ can be written as the gradient of $\phi(\mathbf{p}_{j},\mathbf{o})$ w.r.t. the point coordinate $\mathbf{p}_{j}$. Softmin is also used for differentiability:
@@ V EXPERIMENTS
In this section, we aim to verify: (1) whether finger-rolling behavior is achievable from the optimization, (2) whether finger-rolling and geometry information improve the performance over baseline, (3) whether our method can be applied to a real-world robotic system, and (4) whether our method is fast enough to be applied in an MPC framework.
To answer the questions above, we design two tasks: peg alignment and screwdriver turning. We also test our method on the screwdriver turning task in a real-world setup.
V-A Experiment Setup
In our experiments, we assume two manipulation phases: pregrasping and turning. For pregrasping, we slightly modify our trajectory optimization: we remove all the constraints except the contact constraints at the final step. Since pregrasping and choosing contact points are not the scope of this work, we choose the robot’s initial pose so that the fingers are very close to the desired contact points. This section mainly focuses on the turning phase. During the turning phase, the robot is not allowed to break or establish new contacts, i.e., it is not allowed to change the contact mode.
@@ V-A Experiment Setup
We use the Allegro hand [33] as our robot, which consists of four fingers and sixteen servo-driven joints. Wrist movement is not considered in our experiments. Additionally, we only consider the thumb, index finger, and middle finger, as three fingers are sufficient for all the tasks we consider. Isaac Gym is used as our simulator. [34].
For the trajectory optimization problem setup, we use the horizon $T=12$ steps for the screwdriver turning and $T=10$ steps for peg alignment. The horizon $T$ is reduced by 1 after each step of execution. To solve the trajectory optimization, we use 300 steps at the first step to warm up, then 50 steps for the remaining steps in simulation, and 30 steps in the real world. We reduce the iterations for real-world experiments to slightly speed up the algorithm. The solution of the previous step is used to initialize the solution for the current step. The initialization for actions at the first step is sampled from normal distributions with a mean of 0. The standard deviation for delta actions $\hat{\mathbf{u}}_{t}$ is $2.5\times\mathrm{10}^{-2}$ and for force is between $1.5\times\mathrm{10}^{-2}$ and $1.5$, which is manually tuned depending on the task and which finger to use. The initialization for the joint configurations is produced by iteratively adding the delta actions to the previous joint configurations, i.e., $\mathbf{q}_{t+1}=\mathbf{q}_{t}+\hat{\mathbf{u}}_{t}$. The standard deviation of the normal distribution for each dimension is tuned according to the task. Object configurations $\mathbf{o}_{t}$ are initialized via linear interpolation between the current and goal configurations. SLERP [35] is used to linearly interpolate the orientations.
In our experiments, we consider the following two tasks (See Fig. 5):
Figure 5: The robot needs to turn the screwdriver 90 degrees and turn the peg upright. Each method is visualized with the most common behavior observed in the experiment. The frames shown in the figure are not sampled with the same interval but are manually chosen to better demonstrate the performance of each method. Our method is the only method that can reorient the object very close to the desired pose.
@@
Peg Alignment: Peg-in-hole tasks are important in industrial applications such as robot assembly. Usually, the first step to insert a peg is to reorient and align it with the hole. Reorienting the peg requires finger rolling. Additionally, we include a box to aid with completing the task, e.g., to reorient the peg to an upright pose, one can push the peg to the vertical surface of the box. This behavior is called extrinsic dexterity [36, 37]. This experiment also highlights the versatility of our framework, because including extrinsic contact only requires adding one constraint.
@@
The environment consists of a peg and a box. The peg is a $4~{}\text{cm}\times 4~{}\text{cm}\times 30~{}\text{cm}$ cuboid. The peg starts in contact with the box. The robot needs to utilize the box to turn the peg for approximately 45 degrees to make it upright.
@@
Screwdriver Turning: We assume the screwdriver is already mated with the screw head and focus on turning. The goal is to turn the screwdriver 90 degrees while keeping it upright. The screwdriver turning task can benefit from rich finger-rolling behavior. It also tests the performance of our method on everyday tool manipulation tasks.
[…]
Since the precision screwdriver is designed to be used with an index finger pushing the top, we introduce another constraint making sure that the contact between the index finger and the screwdriver only happens at the desired region (top): $||\mathbf{c}_{0,t}-\hat{\mathbf{c}}(\mathbf{o}_{t})||-h\leq 0,$ where $\mathbf{c}_{0,t}$ is the contact point between the index finger and the object, $h=2~{}\text{cm}$, and $\hat{\mathbf{c}}(\mathbf{o}_{t})$ is the ideal contact location, which is the center of the top surface of the screwdriver.
@@
Distance to Goal: The metric evaluates the distance between the object’s final and desired orientation, defined as the relative angle between two SO(3) orientations.
[…]

### v2

@@ IV-A3 Wrench Balance Constraint
$\sum_{i}\mathbf{f}_{i,t}+\mathbf{f}_{e,t}=m\mathbf{g}$ (4)
$\sum_{i}\mathbf{f}_{i,t}\times\mathbf{r_{i}}(\mathbf{s}_{t})+\mathbf{f}_{e,t}\times\mathbf{r_{e}}(\mathbf{s}_{t})=m\mathbf{r_{com}}(\bm{\theta}_{t+1})\times\mathbf{g}$ (5)
where $K_{p}$ is the gain matrix for the PD controller, $\bm{\tau}_{g}(\mathbf{q}_{t})$ is the joint torque required to counteract the gravity, and $\mathbf{g}=[0,0,9.8]^{T}$ is the gravity vector. $\mathbf{r_{i}}$, $\mathbf{r_{e}}$, and $\mathbf{r_{com}}$ are the vectors represented in the world frame pointing from the object frame origin to the contact point with the $i$th finger, the contact point with the environment and the object’s CoM, respectively. $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ are the force that the $i$th finger and the environment apply to the object respectively. The visualization of the variables is shown in Fig. 2c. Eq. 3, Eq. 4 and Eq. 5 refer to the torque balance for robot joints, the force and torque balance for the object respectively.
Having both control input $\hat{\mathbf{u}}_{t}$ and contact forces $\mathbf{f}_{i,t}$ might seem redundant, since one can possibly derive $\mathbf{f}_{i,t}$ from $\mathbf{s}_{t}$ and $\hat{\mathbf{u}_{t}}$. However, explicitly computing the contact forces is not trivial. Previous work has proposed to include computing contact forces as part of the optimization problem [36, 37]. Similar to those methods, we add the forces $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ as decision variables and compute them within the optimization. The wrench balances also ensure the consistency of control actions $\hat{\mathbf{u}}_{t}$ and forces.
@@ IV-A4 Friction Constraints
We use the Coulomb friction constraint: $||\mathbf{f}_{i,t}^{t}||\leq\mu||\mathbf{f}_{i,t}^{n}||,$ where $\mathbf{f}_{i,t}^{t}$ is the tangential force, $\mathbf{f}_{i,t}^{n}$ is the normal force, and $\mu$ is the friction coefficient. However, satisfying the second-order friction constraint can be numerically unstable when the tangential force is close to 0. To address this, we use a linearized 4-sided friction cone [38], formulated as: $A(\mathbf{n}_{i,t},\mu)\mathbf{f}_{i,t}\leq 0$, where the matrix $A$ depends on the contact normal $\mathbf{n}_{i,t}$ and $\mu$.
@@ IV-A5 Trajectory Optimization
The trajectory optimization problem is:
$\min_{\begin{subarray}{c}\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{T};\\ \mathbf{u}_{0},\mathbf{u}_{1},\cdots,\mathbf{u}_{T-1}\end{subarray}}J_{goal}(\bm{\tau})+J_{smooth}(\bm{\tau})$ (6)
@@
$\sum_{i}\mathbf{f}_{i,t}+\mathbf{f}_{e,t}=m\mathbf{g}$ (11)
$\sum_{i}\mathbf{f}_{i,t}\times\mathbf{r_{i}}(\mathbf{s}_{t})+\mathbf{f}_{e,t}\times\mathbf{r_{e}}(\mathbf{s}_{t})=m\mathbf{r_{com}}(\bm{\theta}_{t+1})\times\mathbf{g}$ (12)
$A(\mathbf{n}_{i}(\mathbf{s}_{t}),\mu)\mathbf{f}_{i,t}\leq 0$ (13)
$f_{i}^{min}-||\mathbf{f}_{i,t}||\leq 0.$ (14)
Eq. 8, 9, and 13 refer to $f_{contact}$, $f_{kinematics}$, and $f_{friction}$, and Eq. 10, 11, and 12 refers to $f_{balance}$. Eq. 14 refers to $f_{min\_f}$, with $f_{i}^{min}$ as the predefined force threshold.
@@ IV-B Geometry Parametrization
As described above, we utilize geometry information to formulate the constraints, such as contact constraints, and kinematics constraints. However, incorporating geometry information, such as meshes, as input is generally intractable. Computing geometry-related gradients for trajectory optimization algorithms adds further complexity. One of our main contributions can be summarized as using a sample-based method combined with softmin to incorporate shape-related constraints and gradients. Those include the distance function $\Phi_{i}(\mathbf{q}_{t},\mathbf{o}_{t})$ for contact constraints, contact points $\mathbf{c}_{i,t}$, the jacobian matrix at the contact point $J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})$, and the contact normal $\mathbf{n}_{i,t}$ for kinematics constraints. $\mathbf{n}_{i,t}$ is also used in formulating the friction constraints. In this section, we will ignore the timestep subscript $t$ for simplicity.
We assume the manipulated object and the robot are described by primitive shapes and triangle meshes respectively. Primitive shapes include boxes, cylinders, spheres, and their combinations. Many objects in real-world applications can be simplified as primitives or their combinations, e.g., a screwdriver as a combination of two cylinders with different radii. Primitives have an analytical formulation of their geometry and SDF, enabling efficient computation of contacts and distances. In the following text, we will utilize the function $\phi_{k}(\mathbf{p}_{j},\mathbf{o})$ to query the SDF of the $k$th primitive of configuration $\mathbf{o}$ at position $\mathbf{p}_{j}$. The subscript $k$ is introduced because, in practice, the object can be a combination of multiple primitives. Thus, the distance between the point $\mathbf{p}_{j}$ and the object can be written as $\phi(\mathbf{p}_{j},\mathbf{o})=\min_{k}\phi_{k}(\mathbf{p}_{j},\mathbf{o})$
@@
With the assumption of the primitive objects, we have access to the SDF of the object: $\phi(\mathbf{p}_{j},\mathbf{o})$. The distance between the $i$th finger and the object can be written as: $\Phi_{i}(\mathbf{q},\mathbf{o})=\min_{j\in P_{i}}\phi(\mathbf{p}_{j},\mathbf{o})$, where $P_{i}$ is the set of points sampled on the surface of the $i$th fingertip. While this provides an accurate estimation with enough sampled points, the minimization operation is not differentiable and also discards information from the other points. Gradients from these points are important for finger rolling as they convey details about the shape of the object, containing information about how to roll the fingers on the surface of the manipulated object based on their geometries. This information becomes critical when precise control is necessary to successfully complete the task. To address the differentiablity problem, we use the softmin to produce a weighted summation over all the points. The weight $h_{j}$ for the $j$th point is given by:
$h_{j}=\frac{\exp(-\delta\phi(\mathbf{p}_{j},\mathbf{o}))}{\sum_{k}\exp(-\delta\phi(\mathbf{p}_{k},\mathbf{o}))},\vspace{-0.2cm}$ (15)
where $\delta$ is the temperature. In practice, we use a high temperature to make the softmin approximation closely match the actual distance. The estimated distance is written as:
$\Phi_{i}(\mathbf{q},\mathbf{o})=\sum_{j\in P_{i}}h_{j}\phi(\mathbf{p}_{j},\mathbf{o}).\vspace{-0.2cm}$ (16)
@@
$\mathbf{c}_{i}=\sum_{j\in P_{i}}h_{j}\mathbf{p}_{j},\quad J_{\mathbf{c}_{i}}(\mathbf{q})=\sum_{j\in P_{i}}h_{j}J_{\mathbf{p}_{j}}(\mathbf{q}).\vspace{-0.2cm}$ (17)
The contact normal $\mathbf{n}_{i}(\mathbf{p},\mathbf{o})$ is estimated as the gradient of $\phi(\mathbf{p}_{j},\mathbf{o})$ w.r.t. $\mathbf{p}_{j}$ using softmin, ensuring differentiability:
@@ V EXPERIMENTS
In this section, we aim to verify: (1) whether finger-rolling behavior is achievable from the optimization, (2) whether finger-rolling and geometry information improve the performance over baseline, (3) whether our method works on real-world systems, and (4) whether our method is fast enough to be applied in an MPC framework.
To answer the questions above, we propose a benchmark for multi-finger dexterous manipulation including five tasks that vary in difficulty. We also test our method and one of the best-performing baselines in a real-world setup.
V-A Benchmark Setup
@@
The task setup for each experiment is shown in Fig. 4. The Allegro hand [39], which consists of four fingers and sixteen servo-driven joints, is used for all the tasks. We only use the thumb, index finger, and middle finger, as three fingers are sufficient for all the tasks we consider. Additionally, we exclude wrist movements, relying solely on finger action. Isaac Gym is used as our simulator [40].
Valve Turning(VT): The object is a cross-shaped valve with one revolute joint, consisting of two cuboids. The goal is to turn the valve by $45^{\circ}$, set relatively close as the robot cannot rotate the valve significantly without changing the contact mode.
Screwdriver Turning(ST): We assume the screwdriver is already mated with the screw head and focus on turning. The goal is to turn the screwdriver 90 degrees while keeping it upright. The screwdriver turning task can benefit from rich finger-rolling behavior. It also tests the performance of our method on everyday tool manipulation tasks.
Simulating the screwdriver turning behavior is challenging due to the complex screw thread geometry and the interaction between the screwdriver tip and the screw. To simplify, we approximate these interactions in the simulator by attaching the screwdriver to a table using a spherical joint, assuming it remains seated in the screw head. We apply damping in the yaw direction to approximate the required turning torque. We model the screwdriver as two cylinders. In hardware experiments, we use an actual screw.
@@
Cuboid Alignment(CA): Peg-in-hole tasks are important in industrial applications such as robot assembly. Usually, the first step to insert a peg is to reorient and align it with the hole. Reorienting the peg requires finger rolling. In this task, we model the peg as a cuboid. Additionally, we include a box to aid with completing the task, e.g., to reorient the cuboid to an upright pose, one can push the cuboid to the vertical surface of the box. This experiment also highlights the versatility of our framework, because including extrinsic contact only requires adding one constraint.
@@
The environment consists of a cuboid(peg) and a box. The cuboid starts in contact with the box. The robot needs to utilize the box to turn the cuboid for approximately 45 degrees to make it upright.
@@
Cuboid Turning(CT): 6D in-hand reorientation is a fundamental task in multi-finger manipulation. We first start with reorienting a cuboid. The robot starts with the thumb and middle finger positioned on the opposite sides of the cuboid to ensure stability. The objective is to turn the cuboid by $60^{\circ}$, reorienting it horizontally. Unlike most RL methods focusing on in-hand reorientation [41, 42], the hand in our setup faces downward instead of upward, making dropping the object more likely.
Complex Reorientation(CR): This task also focuses on 6D in-hand reorientation but is more challenging than the cuboid turning task. The robot needs to reorient a thin and complex object composed of multiple cuboid primitives. This task is designed to evaluate the ability to manipulate complex objects composed of multiple primitives.
In our experiments, we focus exclusively on the stage where the robot has already established contact with the object. Simulations are initialized either with the robot in contact or using a pregrasp algorithm, detailed in Appendix.
For the trajectory optimization problem setup, we use the horizon $T=12$ steps for ST and $T=10$ steps for the rest of the tasks, based on how much the object needs to rotate. The horizon $T$ is reduced by 1 after each step of execution. To solve the trajectory optimization, we use relatively long iterations at the first step to warm up, then much shorter iterations for the remaining steps in the simulation. The solution of the previous step is used to initialize the solution for the current step. The initialization for robot states and actions at the first step is sampled from normal distributions. See details in the appendix. Object configurations $\mathbf{o}_{t}$ are initialized via linear interpolation between the current and goal configurations.
Figure 4: The benchmark consists of multi-finger manipulation tasks with and without extrinsic environment contacts.
Method Valve Turning Screwdriver Turning Cuboid Alignment Cuboid Turning Complex Reorientation
distance valid distance valid distance valid distance valid distance valid
Ours $5.45^{\circ}\pm 2.67^{\circ}$ n.d. $22.84^{\circ}\pm 5.61^{\circ}$ 0.93 $14.84^{\circ}\pm 1.05^{\circ}$ 1.0 $4.57^{\circ}\pm 1.02^{\circ}$ 1.0 $21.78^{\circ}\pm 3.59^{\circ}$ 0.97
Ablation $3.45^{\circ}\pm 1.93^{\circ}$ n.d. $21.16^{\circ}\pm 4.18^{\circ}$ 0.07 $18.66^{\circ}\pm 4.43^{\circ}$ 0.83 $72.12^{\circ}\pm 17.21^{\circ}$ 0.23 $35.04^{\circ}\pm 20.11^{\circ}$ 0.47
PtC $8.02^{\circ}\pm 3.97^{\circ}$ n.d. No valid trial 0.0 No valid trial 0.0 $8.01^{\circ}\pm 3.97^{\circ}$ 1.0 No valid trial 0.0
MPPI $1.77^{\circ}\pm 1.74^{\circ}$ n.d. $33.32^{\circ}\pm 14.61^{\circ}$ 0.83 $5.67^{\circ}\pm 1.75^{\circ}$ 0.87 $8.65^{\circ}\pm 6.26^{\circ}$ 0.80 $28.90^{\circ}\pm 26.78^{\circ}$ 0.53
RL $0.65^{\circ}\pm 0.00^{\circ}$ n.d. No valid trials 0.0 $90.01^{\circ}\pm 0.02^{\circ}$ 1.0 $90.39^{\circ}\pm 1.21^{\circ}$ 1.0 $58.98^{\circ}\pm 34.03^{\circ}$ 1.0
TABLE I: Results are evaluated for 30 episodes in the simulation. N.D. means not defined.
@@
Validity: The algorithm may exploit simulator artifacts to achieve the goal. To address this, we design a validity metric based on human task performance to capture physical plausibility. Specifically, the screwdriver should be upright. A trajectory is considered valid if the center of the top of the screwdriver handle remains within 2cm of its starting position. For the remaining 6D reorientation tasks, validity is defined as not dropping the object. validity is not applicable in the valve turning tasks (marked as n.d. in Tab. I).
Distance to Goal: This metric evaluates the distance between the object’s final and desired orientation, defined as the relative angle between two SO(3) orientations. Only valid trajectories are considered for the evaluation.